A uniform 2.45m ladder of mass 7.70kg is leaning against a vertical wall while making an angle of 53.0? with the floor. A worker pushes the ladder up against the wall until it is vertical.
Answer #1

The center of mass of the vertical ladder is 2.45/2 = 1.225 m above the floor

The center of mass at the 53 degree angle is from the sin function...

sin(53) = y/1.225

y = .978 m

The distance it rises is 1.225 - .978 - .247 m

W = mgh

W = (7.70)(9.8)(.247)

W = 18.6 Joules